Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $111,800 of manufacturing overhead for an estimated allocation base of $86,000 direct material dollars to be used in production. The company has provided the following data for the just completed year:

Purchase of raw materials $ 136,000
Direct labor cost $ 82,000
Manufacturing overhead costs:
Indirect labor $ 115,000
Property taxes $ 8,100
Depreciation of equipment $ 17,000
Maintenance $ 16,000
Insurance $ 10,500
Rent, building $ 36,000
Beginning Ending
Raw Materials $ 27,000 $ 10,000
Work in Process $ 45,000 $ 35,000
Finished Goods $ 71,000 $ 57,000

Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $35,000 ending balance in Work in Process includes $8,800 of direct materials. Given this assumption, supply the information missing below:

  • Required 1
  • Required 2
  • Required 3
  • Required 4
  • Required 5

Compute the predetermined overhead rate for the year.

  • Required 1
  • Required 2
  • Required 3
  • Required 4
  • Required 5

Compute the predetermined overhead rate for the year.

Predetermined overhead rate %

Compute the amount of underapplied or overapplied overhead for the year.

   overhead   
  • Required 3
  • Required 4
  • Required 5

Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

Gitano Products
Schedule of Cost of Goods Manufactured
Direct materials:
Total raw materials available
Raw materials used in production
Total manufacturing costs
  
Cost of goods manufactured   
  • Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

    Unadjusted cost of goods sold =   
  • Assume that the $35,000 ending balance in Work in Process includes $8,800 of direct materials. Given this assumption, supply the information missing below:

    Direct materials   
    Direct labor
    Manufacturing overhead
    Work in process inventory $0
Answer #1

Solution 1:

Predetermined overhead rate = Estimated overhead / estimated direct materials dollars = $111800/86000 = 130%

Solution 2:

Direct materials used in production = beginning materials + purchase of raw materials - ending raw materials = 27000 +136000 -10000 = $153,000

Overhead applied = $153000*130% = $198,900

Actual overhead = 115000+8100+17000+16000+10500+36000 = $202,600

Underapplied overhead = $202600 - $198900 = $3,700

Solution 3:

D E F G 1 2 Gitano Products 5 Schedule of cost of goods manufactured 6 Direct Material: 7 Raw Material, Beginning Add: Purcha

Solution 4:

К 1 2 5 Gitano Products Computation of Unadjusted cost of goods sold 6 Cost of goods manufactured Add: Finished Goods, Beginn

Solution 5:

Direct materials = $8,800

Manufacturing overhead = $8800*130% = $11,440

Direct labor = $35000 - $8800 -$11440 = $14,760

Total work in process inventory = $35,000